Compound Amount


Adah discovered a saving account left to her by a rich uncle while cleaning his old house. When annually if she is now 20 years old.how much is the account worth


1
Expert's answer
2021-03-30T07:19:39-0400

Compound Interest

"A= P(1+r\/n)^{nt}"

Where

A= Total Amount

P= Principal amount

r= annual rate of interest

n= Number of times interest is compounded in a year

t= how long the money is compounded

where n = 12 and t= 20

A= P ( 1+r/12)240

or

A =P (1+i)20